# Deep Unfolding-Aided Parameter Tuning for Plug-and-Play Based Video Snapshot Compressive Imaging

### v2

@@ Abstract
Snapshot compressive imaging (SCI) captures high-dimensional data efficiently by compressing it into two-dimensional observations and reconstructing high-dimensional data from two-dimensional observations with various algorithms. Plug-and-play (PnP) is a promising approach for the video SCI reconstruction because it can leverage both the observation model and denoising methods for videos. This paper proposes a deep unfolding-based method for tuning noise level parameters in PnP-based video SCI, which significantly affects the reconstruction accuracy. For the training of the parameters, we prepare training data from the densely annotated video segmentation (DAVIS) dataset, reparametrize the noise level parameters, and apply the checkpointing technique to reduce the required memory. Simulation results show that the trained noise level parameters significantly improve the reconstruction accuracy and exhibit a non-monotonic pattern, which is different from the assumptions in the conventional convergence analyses of PnP-based algorithms.
@@
Snapshot compressive imaging, deep unfolding, plug and play, parameter tuning.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I INTRODUCTION
Snapshot compressive imaging (SCI) [1] has gained attention as an approach to capturing high-dimensional data. In SCI, high-dimensional data is modulated, compressed into two-dimensional data, and observed with a camera. We then reconstruct the original high-dimensional data using some reconstruction algorithm. Since the data is compressed before observation, it is possible to acquire high-dimensional data with low memory and power consumption. SCI has been used in various fields, including video processing [2, 3, 4, 5], hyperspectral imaging [6, 7, 8, 9, 10, 11], holography [12, 13, 14], depth sensing [15, 16], and polarization imaging [17].
@@
Video SCI is a technique to obtain high frame rate videos by capturing information over multiple time frames. In video SCI, the unknown video of a scene is modulated by a mask and observed as a compressed image. The mask determines which pixels are observed in the compressed image and changes more rapidly than the frame rate of the camera. The high frame rate video is then reconstructed from the measurement image and the information of the mask.
@@
The reconstruction methods for video SCI are mainly categorized into three types: mathematical optimization-based methods [18, 19, 20, 21, 22], machine learning-based methods [23, 24, 25, 26], and hybrid methods [27, 28, 29, 30, 31, 32, 33, 34, 35]. In mathematical optimization-based methods, we first design an optimization problem and then construct an algorithm for the problem. We can incorporate various information on the measurement model and properties of videos directly in the design of the optimization problem. On the other hand, in machine learning-based methods, we construct a machine learning model like a deep neural network and train the model using data of the videos and the corresponding measurements. Such an approach has the potential to reconstruct the unknown videos more accurately with lower computational complexity. For example, reversible SCI net [24] can achieve better reconstruction accuracy than that of a mathematical optimization-based method named Decompress SCI (DeSCI) [21]. However, learning-based methods require the training phase for the training of many parameters. Moreover, these methods often lack interpretability and flexibility for changes in the measurement model compared to mathematical optimization-based methods.
@@
To take advantage of both methods, some hybrid approaches combining mathematical optimization and machine learning have been proposed for video SCI [27, 28, 29, 30, 31, 32, 33, 34, 35]. Most of these methods are based on the plug-and-play (PnP) approach [36, 37], where an update equation of the optimization algorithm is replaced with a denoising method for natural images or videos. For example, on the basis of generalized alternating projection (GAP) [38, 20] and alternating direction method of multipliers (ADMM) [39, 40, 41, 42], PnP-GAP and PnP-ADMM have been proposed for video SCI, respectively [27, 28]. For the incorporated denoising method, various methods such as fast and flexible denoising network (FFDNet) [43] and fast deep video denoising net (FastDVDnet) [44] have been utilized. These PnP methods have lower computational complexity than that of the optimization-based DeSCI, while the learning-based high-quality denoising can be used in the algorithm.
@@
In PnP-based methods, we need to design noise level parameters, i.e., the strength of the denoising at each iteration. The choice of the parameters significantly affects the final reconstruction accuracy. In conventional studies, however, the noise level parameters have been manually designed and have not been discussed deeply. To maximize the potential of PnP-based methods and achieve better reconstruction performance, we need to examine the effect of these parameters in more detail.
@@
In this paper, we propose a deep unfolding-based tuning method for the noise level parameters in the PnP approach. Deep unfolding [45, 46, 47] leverages the similarity between the structure of iterative algorithms and neural networks to train appropriate values for internal algorithm parameters via automatic differentiation. Since the update equations of PnP-GAP and PnP-ADMM are differentiable with respect to the noise level parameters, deep unfolding can be used to train appropriate noise levels in principle. When we use deep learning-based denoising like FastDVDnet, however, we require huge GPU memory to train the noise level parameters. To reduce the required memory, we use the checkpointing technique [48, 49] in the proposed method. In computer simulations, we prepare the data for the training, train the noise level parameters, and evaluate the reconstruction performance with the conventional parameter settings. Simulation results show that the trained parameters via the proposed approach can achieve better reconstruction accuracy. Moreover, the trained parameters exhibit a non-monotonic behavior along the iterations of the reconstruction algorithms. This implies that the monotonically decreasing noise levels assumed for the convergence proof [50, 28, 34, 35] are not necessarily optimal in terms of reconstruction accuracy. This interesting result shows similarity to the trained parameters of some other deep-unfolded algorithms [51, 52]. Our results would also provide insights into the research fields of deep unfolding.
@@
  • In the simulations, we show that the trained noise level parameters have a zigzag behavior, which is similar to those of other deep unfolding approaches [51, 52]. This behavior is different from the usual assumption for the convergence guarantee [50, 28, 34, 35], and hence our results will provide motivation for the convergence analysis of PnP methods with non-decreasing noise level parameters.
The rest of this paper is organized as follows. In Section II, we describe the system model of video SCI. In Section III, we review some PnP-based conventional methods for video SCI. In Section IV, we propose the parameter tuning method based on deep unfolding. In Section V, we present the results of computer simulations. Finally, we conclude the paper in Section VI.
In this paper, we use the following notations. $\mathbb{R}$ denotes the set of all real numbers, and $\mathbb{N}_{+}$ denotes the set of all positive integers. $(\cdot)^{\top}$ and $(\cdot)^{-1}$ represent the transpose and inverse of a matrix, respectively. $\odot$ denotes the Hadamard product. For a vector $\bm{a}=[a_{1},\dotsc,a_{N}]^{\top}\in\mathbb{R}^{N}$, $\left\lVert\bm{a}\right\rVert_{2}=\sqrt{\sum_{n=1}^{N}a_{n}^{2}}$ denotes the $\ell_{2}$ norm of the vector, and $\mathrm{diag}(a_{1},\dotsc,a_{N})\in\mathbb{R}^{N\times N}$ denotes the diagonal matrix whose diagonal elements are $a_{1},\dotsc,a_{N}$. For a matrix $\bm{A}=\left[\bm{a}_{1}\ \dotsb\ \bm{a}_{M}\right]\in\mathbb{R}^{N\times M}$, $\mathrm{vec}(\bm{A})=\left[\bm{a}_{1}^{\top}\ \dotsb\ \bm{a}_{M}^{\top}\right]^{\top}\in\mathbb{R}^{MN}$ denotes the vectorization of $\bm{A}$ obtained by stacking the columns of $\bm{A}$. For a function $f:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$, the proximity operator of $f(\cdot)$ is defined as $\mathrm{prox}_{f}(\bm{s})=\operatornamewithlimits{\mathrm{arg\ min}}_{\bm{x}\in\mathbb{R}^{N}}\left\{f(\bm{x})+\frac{1}{2}\left\lVert\bm{x}-\bm{s}\right\rVert_{2}^{2}\right\}$.
@@ II System Model
Let $\bm{X}\in\mathbb{R}^{N_{x}\times N_{y}\times N_{t}}$ represent an unknown grayscale video of $N_{t}$ frames with image size $N_{x}\times N_{y}$ corresponding to a scene. In SCI systems, the frames of the scene are modulated by a mask $\bm{M}\in\left\{0,1\right\}^{N_{x}\times N_{y}\times N_{t}}$ and observed as a two-dimensional compressed measurement
@@
Here, $\bm{Y}$ is the measurement, $\bm{X}_{n_{t}}=\bm{X}(:,:,n_{t})\in\mathbb{R}^{N_{x}\times N_{y}}$ denotes the $n_{t}$-th frame of the video, $\bm{M}_{n_{t}}=\bm{M}(:,:,n_{t})\in\mathbb{R}^{N_{x}\times N_{y}}$ denotes the corresponding $n_{t}$-th mask, and $\bm{E}\in\mathbb{R}^{N_{x}\times N_{y}}$ denotes the measurement noise. Since $\odot$ denotes the Hadamard product, one in the mask means that the corresponding pixel is observed, and zero means that the pixel is not observed.
@@
$\bm{y}$ $=\bm{\Phi}\bm{x}^{\ast}+\bm{e}.$ (2)
Here, $\bm{y}=\mathrm{vec}(\bm{Y})\in\mathbb{R}^{N_{x}N_{y}}$ and $\bm{e}=\mathrm{vec}(\bm{E})\in\mathbb{R}^{N_{x}N_{y}}$ are the vectorized forms of the measurement image and the noise, respectively. $\bm{\Phi}=[\bm{D}_{1}\ \dotsb\ \bm{D}_{N_{t}}]\in\mathbb{R}^{N_{x}N_{y}\times N_{x}N_{y}N_{t}}$ and $\bm{x}^{\ast}=\left[\bm{x}_{1}^{\top}\ \dotsb\ \bm{x}_{N_{t}}^{\top}\right]^{\top}\in\mathbb{R}^{N_{x}N_{y}N_{t}}$ are the corresponding measurement matrix and the unknown vector, respectively, where $\bm{x}_{n_{t}}=\mathrm{vec}(\bm{X}_{n_{t}})\in\mathbb{R}^{N_{x}N_{y}}$ represents the vectorization of the $n_{t}$-th frame and $\bm{D}_{n_{t}}=\mathrm{diag}(\mathrm{vec}(\bm{M}_{n_{t}}))\in\mathbb{R}^{N_{x}N_{y}\times N_{x}N_{y}}$ represents a diagonal matrix whose diagonal elements are the vectorized forms of $\bm{M}_{n_{t}}$ ($n_{t}=1,\dotsc,N_{t}$). The goal of video SCI is to reconstruct the unknown video $\bm{x}^{\ast}$ from the compressed measurement $\bm{y}$ and the information of the mask $\bm{\Phi}$.
@@ III Conventional Methods
In this section, we briefly review some PnP-based conventional methods for video SCI.
@@ III-A PnP-ADMM
Since the function $R(\cdot)$ is a regularizer for videos, it should return small values for clean videos. The right hand side of (15) means a clean video close to $\bm{x}^{k+1}+\bm{u}^{k}$, which can be interpreted as denoising for the video $\bm{x}^{k+1}+\bm{u}^{k}$. From this viewpoint, PnP approach replaces the update of $\bm{v}^{k}$ with a denoising $\mathcal{D}_{\sigma_{k}}(\cdot)$, i.e.,
@@
Here, $\sigma_{k}\in[0,1]$ is the noise level parameter to determine the strength of denoising. As the denoising function $\mathcal{D}_{\sigma_{k}}(\cdot)$, we can use various denoising methods such as FFDNet [43] and FastDVDnet [44]. In [28], FastDVDnet is used because it is trained for video denoising and can achieve high-quality denoising. Using powerful denoising methods improves the reconstruction accuracy of the algorithm empirically. To obtain good reconstruction accuracy, however, the parameters $\sigma_{k}\in[0,1]$ ($k=0,1,\ldots$) need to be appropriately chosen.
@@ III-C Noise Level Paremeters in PnP Methods
In PnP-ADMM and PnP-GAP, we have parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ to control the strength of denoising at each iteration. The setting of these parameters significantly affects the accuracy of video reconstruction. Although FFDNet and FastDVDnet themselves can use noise level map with the same size as the input image, the noise level parameters in PnP methods are usually treated as scalar values, which means that each element of the noise level map is the same value among all pixels. In the previous study [28], the following three settings are considered for 60 iterations:
@@
    $\sigma_{0}=\dotsb=\sigma_{19}=50/255$, $\sigma_{20}=\dotsb=\sigma_{39}=25/255$, $\sigma_{40}=\dotsb=\sigma_{59}=12/255$
@@
    $\sigma_{0}=50/255,\sigma_{k}=0.97\sigma_{k-1}$ ($k=1,\dotsc,K-1$)
@@
    $\sigma_{k}=12/255$ ($k=0,1,\dotsc,K-1$)
Particularly in cases without preprocessing, it has been found in [28] that the first setting achieves the best reconstruction accuracy in simulations. However, the setting is not necessarily optimal, and we need to tune the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ manually to obtain better reconstruction accuracy.

### v3

# Deep Unfolding-Aided Parameter Tuning for Plug-and-Play-Based Video Snapshot Compressive Imaging

@@ Abstract
Snapshot compressive imaging (SCI) captures high-dimensional data efficiently by compressing it into two-dimensional observations and reconstructing high-dimensional data from two-dimensional observations with various algorithms. The plug-and-play (PnP) method is a promising approach for the video SCI reconstruction because it can leverage both observation models and denoising methods for videos. Since the reconstruction accuracy significantly depends on the choice of noise level parameters, this paper proposes a deep unfolding-based method for tuning these parameters in PnP-based video SCI. For the training of the parameters, we prepare training data from the densely annotated video segmentation dataset, reparametrize the noise level parameters, and apply the checkpointing technique to reduce the required memory. Simulation results show that the trained noise level parameters via the proposed approach exhibit a non-monotonic pattern, which is different from the assumptions in the conventional convergence analyses of PnP-based algorithms. These findings provide new insights into both the application of deep unfolding and the theoretical basis of PnP algorithms.
@@
Deep unfolding, parameter tuning, plug and play, snapshot compressive imaging.
@@ I INTRODUCTION
Snapshot compressive imaging (SCI) [1] has gained attention as an approach to capture high-dimensional data. In SCI, high-dimensional data are modulated, compressed into two-dimensional data, and observed using a camera. We then reconstruct the original high-dimensional data using a reconstruction algorithm. As the data are compressed before observation, it is possible to acquire high-dimensional data with low memory and power consumption. SCI has been used in various fields, including video processing [2, 3, 4, 5], hyperspectral imaging [6, 7, 8, 9, 10, 11], holography [12, 13, 14], depth sensing [15, 16], and polarization imaging [17].
@@
Video SCI is a technique used to obtain high frame rate videos by capturing information over multiple time frames. In video SCI, the unknown video of a scene is modulated by a mask and observed as a compressed image. The mask determines which pixels are observed in the compressed image and changes more rapidly than the frame rate of the camera. The high frame rate video is then reconstructed from the measurement image and the mask.
@@
The reconstruction methods for video SCI are mainly categorized into three types: mathematical optimization-based methods [18, 19, 20, 21, 22], machine learning-based methods [23, 24, 25, 26], and hybrid methods [27, 28, 29, 30, 31, 32, 33, 34, 35]. In mathematical optimization-based methods, we first design an optimization problem and then construct an algorithm for the problem. We can incorporate various types of information on the measurement model and properties of videos directly in the design of the optimization problem. In contrast, in machine learning-based methods, we construct a machine learning model such as a deep neural network and train the model using data of the videos and the corresponding measurements. This approach has the potential to reconstruct unknown videos more accurately with lower computational complexity. For example, reversible SCI net [24] can achieve better reconstruction accuracy than a mathematical optimization-based method called decompress SCI (DeSCI) [21]. However, learning-based methods require the training phase to train many parameters. Moreover, these methods often lack interpretability and flexibility for changes in the measurement model compared with mathematical optimization-based methods.
@@
To take advantage of both methods, hybrid approaches combining mathematical optimization and machine learning have been proposed for video SCI [27, 28, 29, 30, 31, 32, 33, 34, 35]. Most of these methods are based on the plug-and-play (PnP) approach [36, 37], where one of update equations of the optimization algorithm is replaced with a denoising method for natural images or videos. For example, on the basis of generalized alternating projection (GAP) [38, 20] and the alternating direction method of multipliers (ADMM) [39, 40, 41, 42], PnP-GAP and PnP-ADMM have been proposed for video SCI [27, 28]. For the incorporated denoising method, various methods such as fast and flexible denoising network (FFDNet) [43] and fast deep video denoising net (FastDVDnet) [44] have been utilized. These PnP methods have lower computational complexity than optimization-based DeSCI, whereas the learning-based high-quality denoising can be used in the algorithm.
@@
In PnP-based methods, we need to design noise level parameters, i.e., the strength of denoising at each iteration. The choice of the parameters significantly affects the final reconstruction accuracy. In conventional studies, however, the noise level parameters have been manually designed and have not been discussed deeply. To maximize the potential of PnP-based methods and achieve better reconstruction performance, we need to examine the effect of these parameters in more detail.
@@
In this paper, we propose a deep unfolding-based tuning method for the noise level parameters in the PnP approach. Deep unfolding [45, 46, 47] leverages the similarity between the structure of iterative algorithms and neural networks to train appropriate values for internal algorithm parameters via automatic differentiation. Since the update equations of PnP-GAP and PnP-ADMM are differentiable with respect to the noise level parameters, deep unfolding can be used to train appropriate noise levels in principle. When we use deep learning-based denoising such as FastDVDnet, however, we require huge GPU memory to train the noise level parameters. To reduce the required memory, we use the checkpointing technique [48, 49] in the proposed method. In computer simulations, we prepare the data for the training, train the noise level parameters, and evaluate the reconstruction performance using the conventional parameter settings. Simulation results show that the parameters trained via the proposed approach can achieve better reconstruction accuracy. Moreover, the trained parameters exhibit non-monotonic behavior during the iterations of the reconstruction algorithms. This implies that the monotonically decreasing noise levels assumed for the convergence proof [50, 28, 34, 35] are not necessarily optimal in terms of reconstruction accuracy. This interesting result is similar to the trained parameters of some other deep-unfolded algorithms [51, 52]. Our results would also provide insights into the research field of deep unfolding.
@@
  • In the simulations, we show that the trained noise level parameters exhibit zigzag behavior, which is similar to those of other deep unfolding approaches [51, 52]. This behavior is different from the usual assumption for the convergence guarantee [50, 28, 34, 35]; hence, our results provide motivation for the convergence analysis of PnP methods with non-decreasing noise level parameters.
The rest of this paper is organized as follows. In Section II, we describe the system model of video SCI. In Section III, we review some conventional PnP-based methods for video SCI. In Section IV, we propose a parameter tuning method based on deep unfolding. In Section V, we present the results of the computer simulations. Finally, we conclude the paper in Section VI.
In this paper, we use the following notation: $\mathbb{R}$ denotes the set of all real numbers, and $\mathbb{N}_{+}$ denotes the set of all positive integers. $(\cdot)^{\top}$ and $(\cdot)^{-1}$ represent the transpose and inverse of matrix, respectively. $\odot$ denotes the Hadamard product. For a vector $\bm{a}=[a_{1},\dotsc,a_{N}]^{\top}\in\mathbb{R}^{N}$, $\left\lVert\bm{a}\right\rVert_{2}=\sqrt{\sum_{n=1}^{N}a_{n}^{2}}$ denotes the $\ell_{2}$ norm of the vector, and $\mathrm{diag}(a_{1},\dotsc,a_{N})\in\mathbb{R}^{N\times N}$ denotes the diagonal matrix whose diagonal elements are $a_{1},\dotsc,a_{N}$. For a matrix $\bm{A}=\left[\bm{a}_{1}\ \dotsb\ \bm{a}_{M}\right]\in\mathbb{R}^{N\times M}$, $\mathrm{vec}(\bm{A})=\left[\bm{a}_{1}^{\top}\ \dotsb\ \bm{a}_{M}^{\top}\right]^{\top}\in\mathbb{R}^{MN}$ denotes the vectorization of $\bm{A}$ obtained by stacking the columns of $\bm{A}$. For a function $f:\mathbb{R}^{N}\to\mathbb{R}\cup\{+\infty\}$, the proximity operator of $f(\cdot)$ is defined as $\mathrm{prox}_{f}(\bm{s})=\operatornamewithlimits{\mathrm{arg\ min}}_{\bm{x}\in\mathbb{R}^{N}}\left\{f(\bm{x})+\frac{1}{2}\left\lVert\bm{x}-\bm{s}\right\rVert_{2}^{2}\right\}$.
@@ II System Model
Let $\bm{X}\in\mathbb{R}^{N_{x}\times N_{y}\times N_{t}}$ represent an unknown grayscale video of $N_{t}$ frames, with an image size $N_{x}\times N_{y}$ corresponding to a scene. In SCI systems, the frames of the scene are modulated by a mask $\bm{M}\in\left\{0,1\right\}^{N_{x}\times N_{y}\times N_{t}}$ and are observed as a two-dimensional compressed measurement
@@
Here, $\bm{Y}$ is the measurement, $\bm{X}_{n_{t}}=\bm{X}(:,:,n_{t})\in\mathbb{R}^{N_{x}\times N_{y}}$ denotes the $n_{t}$-th frame of the video, $\bm{M}_{n_{t}}=\bm{M}(:,:,n_{t})\in\mathbb{R}^{N_{x}\times N_{y}}$ denotes the corresponding $n_{t}$-th mask, and $\bm{E}\in\mathbb{R}^{N_{x}\times N_{y}}$ denotes measurement noise. Since $\odot$ denotes the Hadamard product, one in the mask means that the corresponding pixel is observed, and zero means that the pixel is not observed.
@@
$\bm{y}$ $=\bm{\Phi}\bm{x}^{\ast}+\bm{e},$ (2)
where $\bm{y}=\mathrm{vec}(\bm{Y})\in\mathbb{R}^{N_{x}N_{y}}$ and $\bm{e}=\mathrm{vec}(\bm{E})\in\mathbb{R}^{N_{x}N_{y}}$ are the vectorized forms of the measurement image and noise, respectively. $\bm{\Phi}=[\bm{D}_{1}\ \dotsb\ \bm{D}_{N_{t}}]\in\mathbb{R}^{N_{x}N_{y}\times N_{x}N_{y}N_{t}}$ and $\bm{x}^{\ast}=\left[\bm{x}_{1}^{\top}\ \dotsb\ \bm{x}_{N_{t}}^{\top}\right]^{\top}\in\mathbb{R}^{N_{x}N_{y}N_{t}}$ are the corresponding measurement matrix and the unknown vector, respectively, where $\bm{x}_{n_{t}}=\mathrm{vec}(\bm{X}_{n_{t}})\in\mathbb{R}^{N_{x}N_{y}}$ represents the vectorization of the $n_{t}$-th frame and $\bm{D}_{n_{t}}=\mathrm{diag}(\mathrm{vec}(\bm{M}_{n_{t}}))\in\mathbb{R}^{N_{x}N_{y}\times N_{x}N_{y}}$ represents a diagonal matrix whose diagonal elements are the vectorized forms of $\bm{M}_{n_{t}}$ ($n_{t}=1,\dotsc,N_{t}$). The goal of video SCI is to reconstruct the unknown video $\bm{x}^{\ast}$ from the compressed measurement $\bm{y}$ and the information of the mask $\bm{\Phi}$.
@@ III Conventional Methods
In TABLE I, we summarize related work and recent advances in video SCI. In this section, we briefly review some conventional PnP-based methods that are closely related to the proposed approach.
TABLE I: Summary of related work and recent advances in video SCI.
Reference Main Approach Summary
J. Yang et al., 2014 [19] Model-based Gaussian mixture model-based SCI for video reconstruction.
X. Yuan et al., 2016 [20] Model-based GAP algorithm for SCI.
Y. Liu et al., 2019 [21] Model-based Rank minimization-based approach for SCI reconstruction.
P. Yang et al., 2020 [22] Model-based SCI using Shearlet transform for improved sparsity representation.
M. Qiao et al., 2020 [23] Machine Learning Deep learning-based SCI reconstruction with convolutional architectures.
Z. Cheng et al., 2021 [24] Machine Learning Memory-efficient reversible SCI Net for large-scale video reconstruction.
Z. Cheng et al., 2023 [25] Machine Learning Recurrent neural network (RNN)-based SCI for efficient video reconstruction.
Y.-C. Miao et al., 2024 [26] Machine Learning Domain factorization-based learning for SCI.
X. Yuan et al., 2022 [28] Hybrid PnP-GAP and PnP-ADMM algorithms for video SCI.
Z. Wu et al., 2023 [29] Hybrid Adaptive parameter tuning for PnP methods to reduce the mismatch between training and test data.
Z. Meng et al., 2023 [30] Hybrid Denoising network training via deep unfolding for SCI.
X. Liu et al., 2023 [31] Hybrid Multi-stage fusion network for SCI with GAP framework.
Y. Zhao et al., 2023 [32] Hybrid Memory-efficient SCI reconstruction using deep equilibrium models.
B. Shi et al., 2023 [34] Hybrid Provable denoising algorithms for SCI with convergence guarantees.
B. Shi et al., 2024 [35] Hybrid Spatial-temporal video denoising methods for SCI with provable guarantees.
@@ III-A PnP-ADMM
Since function $R(\cdot)$ is a regularizer for videos, it should return small values for clean videos. The right hand side of (15) indicates a clean video close to $\bm{x}^{k+1}+\bm{u}^{k}$, which can be interpreted as denoising for video $\bm{x}^{k+1}+\bm{u}^{k}$. From this viewpoint, the PnP approach replaces the update of $\bm{v}^{k}$ with denoising $\mathcal{D}_{\sigma_{k}}(\cdot)$, that is,
@@
Here, $\sigma_{k}\in[0,1]$ is the noise level parameter used to determine the strength of denoising. As the denoising function $\mathcal{D}_{\sigma_{k}}(\cdot)$, we can use various denoising methods such as FFDNet [43] and FastDVDnet [44]. In [28], FastDVDnet is used because it is trained for video denoising and can achieve high-quality denoising. The use of powerful denoising methods improves the reconstruction accuracy of the algorithm empirically. To obtain good reconstruction accuracy, however, the parameters $\sigma_{k}\in[0,1]$ ($k=0,1,\ldots$) need to be appropriately chosen.
@@ III-C Noise Level Paremeters in PnP Methods
In PnP-ADMM and PnP-GAP, we have the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ to control the strength of denoising at each iteration. The setting of these parameters significantly affects the accuracy of the video reconstruction. Although FFDNet and FastDVDnet can use a noise level map of the same size as the input image, the noise level parameters in PnP methods are usually treated as scalar values, which means that each element of the noise level map has the same value among all pixels. In the previous study [28], the following three settings are considered for 60 iterations:
@@
    step: $\sigma_{0}=\dotsb=\sigma_{19}=50/255$, $\sigma_{20}=\dotsb=\sigma_{39}=25/255$, $\sigma_{40}=\dotsb=\sigma_{59}=12/255$
@@
    exponential: $\sigma_{0}=50/255,\sigma_{k}=0.97\sigma_{k-1}$ ($k=1,\dotsc,K-1$)
@@
    constant: $\sigma_{k}=12/255$ ($k=0,1,\dotsc,K-1$)
Particularly in cases without preprocessing, it has been found in [28] that the first setting (step) achieves the best reconstruction accuracy in simulations. However, the setting is not necessarily optimal, and we need to tune the parameters $\left\{\sigma_{k}\right\}_{k=0,\dotsc,K-1}$ manually to obtain better reconstruction accuracy.